OpenStudy (anonymous):

Solve for x.Check your answers. A) 20=9+x

OpenStudy (anonymous):

20 = 9 + x (subtract 9 from both sides) 20 - 9 = 9 - 9 + x 11 = x check... 20 = 9 + x 20 = 9 + 11 20 = 20 (correct)